OpenStudy (el_tucan):

find all values in [0,pi) so that dy/dx = 0 for the function y = sin(x^2)

myininaya (myininaya):

Are you having issues finding y'?

myininaya (myininaya):

Do you know the chain rule?

OpenStudy (anonymous):

Ma'am will explain you, don't worry..

OpenStudy (el_tucan):

yeah i get using the chain rule to find the derivative but i don't know what is meant by "finding values so that dy/dx = 0"

myininaya (myininaya):

y'=f(x) y'=0 implies you set f(x)=0 and solve for x

myininaya (myininaya):

Just set the derivative equal to zero and solve for x

myininaya (myininaya):

on the interval [0,pi)

OpenStudy (el_tucan):

i have y' = 2x * cos(x^2) correct? so 2x * cos(x^2) = 0?

myininaya (myininaya):

What did you get for y'?

myininaya (myininaya):

Right! :)

myininaya (myininaya):

So we have x=0 or cos(x^2)=0

myininaya (myininaya):

You need to solve cos(x^2)=0 for x on [0,pi)

OpenStudy (el_tucan):

i plugged cos(0^2) in my graphing calculator which is 1, is that right? im sorry :(

OpenStudy (anonymous):

You are to find values for which it becomes 0 not 1..

myininaya (myininaya):

For what values satisfy cos(u)=0 look at unit circle.

OpenStudy (el_tucan):

gotcha, thanks :)

OpenStudy (anonymous):

Really??

myininaya (myininaya):

You are actually solving cos(u)=0 for u on 0<u<=pi^2 Keep in mind that pi^2 is between 3pi and 7 pi/2 So you actually want to find values for u that satisfy cos(u)=0 on 0<u<3pi

OpenStudy (el_tucan):

sorry waterineyes ;) i got it, thank you myininaya, the square roots of pi/2, 8pi/2, & 5pi/2 are the zero's within that interval

OpenStudy (anonymous):

Sorry for what dear?? I was just confirming that...

OpenStudy (el_tucan):

i thought you were picking on me :) im working on my math esteem lol

OpenStudy (anonymous):

There is nothing like that, I was just confirming that you got it or not...

myininaya (myininaya):

Hey @El_Tucan that 8pi/2 should be something else

myininaya (myininaya):

Those other two values you got were right with the little square root thingy that you said :)